OpenStudy (anonymous):

Factor x^2 - 3 x^2-3^2= (x+3)(x-3) Did I do this right?

OpenStudy (skullpatrol):

nope

OpenStudy (anonymous):

What did I do wrong?

OpenStudy (anonymous):

\[x^2-3=\left( x+\sqrt{3} \right)\left( x-\sqrt{3} \right)\]

OpenStudy (skullpatrol):

Factor x^2 - 9

OpenStudy (anonymous):

I know I was supposed to do the square root thing. But I didn't know how to do it on here.

OpenStudy (anonymous):

How do you factor x^2-9?

OpenStudy (skullpatrol):

x^2-3^2= (x+3)(x-3)

OpenStudy (anonymous):

for this you are correct.

OpenStudy (anonymous):

Okay now I see were I made the mistake. Thank you @skullpatrol and @surjithayer.

OpenStudy (skullpatrol):

Thanks for asking :)